# Stability of Bose-Fermi mixtures in two dimensions: a lowest-order constrained variational approach

### v1

@@
Stability of Bose-Fermi mixtures in two dimensions: a lowest-order constrained variational approach
Pietro Cordioli Dipartimento di Fisica e Astronomia “Augusto Righi”, Università di Bologna, Via Irnerio 46, I-40126, Bologna, Italy    Leonardo Pisani leonardo.pisani2@unibo.it Dipartimento di Fisica e Astronomia “Augusto Righi”, Università di Bologna, Via Irnerio 46, I-40126, Bologna, Italy INFN, Sezione di Bologna, Viale Berti Pichat 6/2, I-40127, Bologna, Italy    Pierbiagio Pieri pierbiagio.pieri@unibo.it Dipartimento di Fisica e Astronomia “Augusto Righi”, Università di Bologna, Via Irnerio 46, I-40126, Bologna, Italy INFN, Sezione di Bologna, Viale Berti Pichat 6/2, I-40127, Bologna, Italy
(17 lutego 2026)
Streszczenie
@@ III Solution of the LOCV equation
Rysunek 1: Dimensionless healing distance ${k_{\mathrm{F}}}d^{-}$ and ${k_{\mathrm{F}}}d^{+}$ versus the coupling strength $\eta=-\ln({k_{\mathrm{F}}}a_{\mathrm{BF}})$ for the attractive and repulsive branches, respectively.
@@
Rysunek 2: (a) Correlation function $f^{-}(r)$ for the attractive branch, for several interaction strengths ranging from $\eta=-2$ (weak attraction) to $\eta=2$ (strong attraction). Inset: pair correlation function $g^{-}(r)=f^{-}(r)^{2}$ for $\eta=-2$ (solid line) together with its small $r$ asymptotic behavior $\propto\log^{2}(r/a_{\mathrm{BF}})$ (dashed-dotted line). (b) Correlation function $f^{+}(r)$ for the repulsive branch, for three interaction strengths ranging from $\eta=0$ (strong repulsion) to $\eta=2$ (weak repulsion). Inset: pair correlation function $g^{+}(r)=f^{+}(r)^{2}$ for $\eta=1$.
@@ IV Attractive and repulsive Energy-Branches
Rysunek 3: Dimensionless quantities $A^{-}$ and $A^{+}$ determining the BF interaction energy as a function of the interaction strength $\eta=-\ln({k_{\mathrm{F}}}a_{\mathrm{BF}})$ for the attractive and repulsive branches, respectively. Insets: (a) comparison between $A^{-}$ (line) and experimental data for the attractive polaron energy normalized by $\epsilon_{\rm F}$ [69] (squares); (b) comparison between $A^{+}$ (line) and QMC results for the polaron energy normalized by $\epsilon_{\rm F}$ for hard-disk BF repulsive interaction [97] (squares); the value obtained in [86] within a $T$-matrix approximation is also reported (asterisk). Dotted lines in both panels: weak-coupling approximation (19); dash-dotted line in (a): strong-coupling approximation (20).
@@
Rysunek 4: Relative quantum depletion coefficients $D^{-}$ (a) and $D^{+}$ (b) as a function of the interaction strength $\eta=-\ln({k_{\mathrm{F}}}a_{\mathrm{BF}})$ for the attractive and repulsive branches, respectively.
@@ V Mechanical Stability
Rysunek 5: Critical value of the BB repulsion strength $\zeta_{c}$ as a function of the BF interaction strength $\eta=-\ln({k_{\mathrm{F}}}a_{\mathrm{BF}})$ for $m_{\rm B}=m_{\rm F}$ and different values of the boson concentration $x$ for the attractive (a) and repulsive (b) branches.
@@
Rysunek 6: Critical BB coupling $\zeta_{\mathrm{c}}$ as a function of the mass ratio $m_{\rm B}/m_{\rm F}$ for $n_{\rm B}=n_{\rm F}$ and different values of the BF coupling strength $\eta$ for the attractive (a) and repulsive (b) branches.
@@ Acknowledgements.
Rysunek 7: (a) Fermionic chemical potential $\mu_{\mathrm{F}}$} and (b) shifted bosonic chemical potential $({\mu_{\mathrm{B}}}+{\varepsilon_{0}})$ (normalized by $\epsilon_{\mathrm{F}}$}) for the attractive branch as a function of the coupling strength $\eta$ for several concentrations $x$. (c) Fermionic chemical potential $\mu_{\mathrm{F}}$} and (d) bosonic chemical potential $\mu_{\mathrm{B}}$} (normalized by $\epsilon_{\mathrm{F}}$}) for the repulsive branch as a function of the coupling strength $\eta$ for several concentrations $x$. In all panels we have used $m_{\rm B}=m_{\rm F}$ and $\zeta=0.1$.
Dodatek A Chemical Potentials and molecular limit
@@
valid to leading order in the small parameters $1/\eta$ and $\zeta\penalty 10000\$[37]. The asymptotic expressions (38) and (39) (with $\epsilon_{0}/{\epsilon_{\mathrm{F}}}$ negligible in their regime of validity) are reported for $x=1$ as black solid lines in panels (a) and (b) of Fig. 7.
@@
Dodatek B Auxiliary conditions to the LOCV Equation

### v2

@@
Abstract
@@ III Solution of the LOCV equation
Figure 1: Dimensionless healing distance ${k_{\mathrm{F}}}d^{-}$ and ${k_{\mathrm{F}}}d^{+}$ versus the coupling strength $\eta=-\ln({k_{\mathrm{F}}}a_{\mathrm{BF}})$ for the attractive and repulsive branches, respectively.
@@
Figure 2: (a) Correlation function $f^{-}(r)$ for the attractive branch, for several interaction strengths ranging from $\eta=-2$ (weak attraction) to $\eta=2$ (strong attraction). Inset: pair correlation function $g^{-}(r)=f^{-}(r)^{2}$ for $\eta=-2$ (solid line) together with its small $r$ asymptotic behavior $\propto\log^{2}(r/a_{\mathrm{BF}})$ (dashed-dotted line). (b) Correlation function $f^{+}(r)$ for the repulsive branch, for three interaction strengths ranging from $\eta=0$ (strong repulsion) to $\eta=2$ (weak repulsion). Inset: pair correlation function $g^{+}(r)=f^{+}(r)^{2}$ for $\eta=1$.
@@ IV Attractive and repulsive Energy-Branches
Figure 3: Dimensionless quantities $A^{-}$ and $A^{+}$ determining the BF interaction energy as a function of the interaction strength $\eta=-\ln({k_{\mathrm{F}}}a_{\mathrm{BF}})$ for the attractive and repulsive branches, respectively. Insets: (a) comparison between $A^{-}$ (line) and experimental data for the attractive polaron energy normalized by $\epsilon_{\rm F}$ [69] (squares); (b) comparison between $A^{+}$ (line) and QMC results for the polaron energy normalized by $\epsilon_{\rm F}$ for hard-disk BF repulsive interaction [97] (squares); the value obtained in [86] within a $T$-matrix approximation is also reported (asterisk). Dotted lines in both panels: weak-coupling approximation (19); dash-dotted line in (a): strong-coupling approximation (20).
@@
Figure 4: Relative quantum depletion coefficients $D^{-}$ (a) and $D^{+}$ (b) as a function of the interaction strength $\eta=-\ln({k_{\mathrm{F}}}a_{\mathrm{BF}})$ for the attractive and repulsive branches, respectively.
@@ V Mechanical Stability
Figure 5: Critical value of the BB repulsion strength $\zeta_{c}$ as a function of the BF interaction strength $\eta=-\ln({k_{\mathrm{F}}}a_{\mathrm{BF}})$ for $m_{\rm B}=m_{\rm F}$ and different values of the boson concentration $x$ for the attractive (a) and repulsive (b) branches.
@@
Figure 6: Critical BB coupling $\zeta_{\mathrm{c}}$ as a function of the mass ratio $m_{\rm B}/m_{\rm F}$ for $n_{\rm B}=n_{\rm F}$ and different values of the BF coupling strength $\eta$ for the attractive (a) and repulsive (b) branches.
@@ Acknowledgements.
Figure 7: (a) Fermionic chemical potential $\mu_{\mathrm{F}}$} and (b) shifted bosonic chemical potential $({\mu_{\mathrm{B}}}+{\varepsilon_{0}})$ (normalized by $\epsilon_{\mathrm{F}}$}) for the attractive branch as a function of the coupling strength $\eta$ for several concentrations $x$. (c) Fermionic chemical potential $\mu_{\mathrm{F}}$} and (d) bosonic chemical potential $\mu_{\mathrm{B}}$} (normalized by $\epsilon_{\mathrm{F}}$}) for the repulsive branch as a function of the coupling strength $\eta$ for several concentrations $x$. In all panels we have used $m_{\rm B}=m_{\rm F}$ and $\zeta=0.1$.
Appendix A Chemical Potentials and molecular limit
@@
valid to leading order in the small parameters $1/\eta$ and $\zeta$~[37]. The asymptotic expressions (38) and (39) (with $\epsilon_{0}/{\epsilon_{\mathrm{F}}}$ negligible in their regime of validity) are reported for $x=1$ as black solid lines in panels (a) and (b) of Fig. 7.
@@
Appendix B Auxiliary conditions to the LOCV Equation